### v2

# thanks: 0009-0007-0442-5443thanks: 0000-0002-4641-022Xthanks: 0000-0001-9980-0363

Abstract
Originally, the Biswas–Chatterjee–Sen model exhibits an order/disorder phase transition for a sufficiently large number of negative interactions among actors. In this paper, the model is extended by the nonconformists and inflexibles. Nonconformists are actors who do not follow the original model rules, but in different ways do something opposite. We introduce inflexibles as actors who does not change their opinions. Both discrete and continuous opinions are considered. With direct Monte Carlo simulations and mean-field calculations, we check the influence of fractions of nonconformists and inflexibles on mean opinion in the system. With the mean-field calculations we identify ranges of fractions of nonconformists where ordered phase of the system is available. The results of the mean-field calculations perfectly match the results of the Monte Carlo simulations. We consider inflexibles adhered: (i) to extreme opinions; (ii) to specific opinions and (iii) chosen independently of their initial opinion. For inflexibles adhered to specific and extreme opinions they play a role of effective bias suppressing disorder phase in the system. The qualitative results of introducing nonconformists (inflexibles) in various ways (discrete/continuous opinions and annealed/quenched disorder) are roughly the same. However, for the model extended by inflexibles, we can observe a systematic shift of the mean order parameter to its higher values for quenched disorder compared with annealed disorder.
I Introduction
Sociophysics [1, 2, 3, 4] tackles the problems of social science with methods and tools of statistical physics. Among these problems, the formation and dynamics of public opinion [5, 6, 7, 8] is one of the core topics of sociophysics.
The opinion formation models may be classified into discrete or continuous ones. The latter are represented by the models of Weisbuch et al. [9, 10] or Hegselmann and Krause [11] applied also to complex networks [12, 13]. In such models, opinions are represented as real numbers, for example, taking values in the interval [1;+1][-1;+1]. Among discrete models one should mention: voter [14, 15, 16, 17], majority-rule [18], Sznajd [19] or models based on social impact [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In such models, the opinions are represented as integers, in the simplest case with only two opinions in $\{-1,+1\}$. However, for discrete models, also more than two opinion values (but still with their finite number) were investigated for the voter model [30, 31, 32, 33, 34, 35, 36, 29], the Sznajd model [37, 38, 39, 29], the Latané model [29, 28, 27, 25, 24], the majority-rule model [40, 41, 42, 43, 44] and in other studies [45, 46, 47, 48, 49].
The kinetic model of opinion dynamics proposed by Biswas, Chatterjee and Sen (BChS) [50, 51] was originally formulated in both versions: continuous and discrete. Soon after original introducing, the BChS model was followed by several modifications [52, 53] with the most recent presented in References 54, 55, 56, 57 (see also References 58, 59 for reviews).
In the BChS model, the interactions can take positive or negative values. In this paper, we enrich the BChS model with inflexibles and nonconformists. Nonconformists are agents who go against the opinion they are supposed to take, and inflexibles are agents who never change their own opinion, but can influence others.
In this context, it can be mentioned that agents having negative interaction in the BChS model have been interpreted as contrarians in a previous study [52], however, we introduce nonconformists as an independent source of noise, irrespective of the type of interaction they have.
Using both mean-field theory calculations and direct Monte Carlo simulation results, we show how the fraction of nonconformists and the fraction of inflexibles influence the original model results. We consider both, discrete and continuous spectrum of opinions. Also various variants of disorder (annealed or quenched) are considered. For annealed disorder, every actor decides to behave as nonconformists (or inflexibles) with a given probability corresponding to the assumed fraction of them in the society. For quenched disorder, the positions of inflexibles and nonconformists are fixed at the network and their fraction are the model control parameters.
Some ways of introduction of inflexibles may shift the original curve of the order parameter vs. noise level towards higher values of the order parameter. As a result, the disordered phase is removed from the system.
In contrast, introducing nonconformists leads to the reduction of the order parameter. For some range of their abundance, only a disordered phase is observed. The results of performed mean-filed calculations perfectly match the results of direct Monte Carlo simulations.
The paper is organized as follows: the Section˜II introduces the model and necessary modifications to deal with groups of inflexibles and nonconformists; Section˜III presents the results (including mean-field calculations in Sections˜III.1, III.2, III.3 and III.4 and Monte Carlo simulations in Section˜III.5) and Section˜IV is devoted to their discussion.
II Model
In the original model formulation [50] the time evolution of the opinion $o_{i}$ of the actor $i$ is given as
$o_{i}(t+1)=o_{i}(t)+\mu_{ij}o_{j}(t),$ (1)
where the elements of the matrix $\mu_{ij}$ take random values from [1;+1][-1;+1] (in continuous version) or $\mu=\pm 1$ (in discrete version). The fraction $p$ of these values is negative and $p$ plays a role of the noise level.
In the continuous version of the model opinions $o_{i}\in[-1;+1]$, while in the discrete version only three values of the opinions are available $o_{i}\in\{-1,0,+1\}$. Anyway, in both versions, if the anticipated value of the opinion according to Equation˜1 is greater (less) than +1+1 (1-1) then it is set to +1+1 (1-1). The latter leads to the rule
$o_{i}(t+1)=\operatorname{\mathrm{sgn}}(o_{i}(t)+\mu_{ij}o_{j}(t))$ (2)
in the discrete version of the model.
The order parameter of the model is defined as
$O(t)=\dfrac{1}{N}\left|\sum_{i}o_{i}(t)\right|.$ (3)
The results of the modeling opinion with the original BChS model [50] can be summarized as follows: the mean-field (fully connected) model shows a order-disorder phase transition as $p$ exceeds a critical value $p_{c}$. $p_{c}$ depends on whether the $\mu$ values are chosen randomly from a discrete set (1-1 or +1+1) or continuously and uniformly between 1-1 to +1+1. $\mu$ can be also considered as either a quenched or annealed variable which does not affect the results. The fully connected model was found to belong to the Ising mean-field universality class. Later, many variations of the original model had been studied [58].
II.1 Nonconformists
According to the Cambridge Dictionary [60], nonconformist is ‘someone who lives and thinks in a way that is different from other people’. This difference from other people may be introduced in various ways.
In the first version, we propose that the fraction $c$ of the actors $i$ applies
$o_{i}(t+1)=o_{i}(t)-\mu_{ij}o_{j}(t)$ (4)
instead of Equation˜1. That is, they subtract the term $\mu_{ij}o_{j}(t)$ instead of adding it.
In the second variant, the fraction $c$ of actors $i$ applies
$o_{i}(t+1)=-[o_{i}(t)+\mu_{ij}o_{j}(t)]$ (5)
instead of Equation˜1. That is, they take an opposite opinion that comes from the original formulation of the BChS model.
In both cases, the remaining part of the society [i.e., fraction (1c)(1-c) of actors] follows Equation˜1.
II.2 Inflexibles
Inflexibles are actors who have fixed opinions and are unable (or unwilling) to change them. For modeling the presence of inflexibles in the society, we assume several scenarios. In the most extreme assumption: the fraction $z$ of the actors adheres to an extreme opinion (either +1+1 or 1-1) and never changes it. We also consider the case where the fraction $z$ of inflexibles is mixed with the believers for both extreme opinions (+1+1 and 1-1). Finally, also the scenario where actors with neutral opinions are inflexibles is considered. The remaining (flexible) part of the population [i.e., the fraction (1z)(1-z) of actors] follows the rule given in Equation˜1.
III Results
III.1 Mean-field calculations
In the mean-field description of the kinetic exchange opinion model and discrete opinions, the state of the system is characterized by the fractions $f_{+}(t)$, $f_{-}(t)$, $f_{0}(t)$ denoting, respectively, the densities of agents holding opinions +1+1, 1-1 and 0. These satisfy the normalized condition
$f_{+}(t)+f_{-}(t)+f_{0}(t)=1.$ (6)
It is convenient to introduce the order parameter
$O(t)=f_{+}(t)-f_{-}(t)$ (7)
and the activity or the non-neutral fraction
$s(t)=f_{+}(t)+f_{-}(t).$ (8)
Thus, $s\in[0;1]$ measures how many agents hold a non-neutral opinion, while $O$ captures net polarization.
III.2 Mean-field calculations for nonconformist (4)
We now introduce nonconformists in the first version according to Equation˜4 following the simplest rule: In an update, the nonconformist agents flips the sign of the influence, i.e., $\mu_{ij}\to-\mu_{ij}$ with probability $c$ and during a normal update with probability 1c1-c, the dynamics remains unchanged.
Consequently, the effective probability that an influence is negative becomes the mixture
$p_{\text{eff}}=(1-c)p+c(1-p)=(1-2c)p+c.$ (9)
In the original model, all mean-field transition rates depend on $\mu$ only through the probability that the influence is negative, namely $p$. In the presence of flip-$\mu$ nonconformists, a fraction $c$ of the updates uses the opposite sign, so the probability that $\mu$ is negative becomes $p_{\text{eff}}$. Therefore, the mean-field evolution in the presence of nonconformists is obtained exactly by replacing $p\to p_{\text{eff}}$. This observation allows us to use the known mean-field equations of the original model while incorporating nonconformists through the single parameter $p_{\text{eff}}$.
Using the compact form of the original mean-field equations and with the replacement $p\to p_{\text{eff}}$ we obtain
$\frac{dO(t)}{dt}=\left[(1-2p_{\text{eff}})-(1-p_{\text{eff}})s\right]O(t)$ (10)
and
$\frac{ds(t)}{dt}=-\frac{1}{2}\left[(2p_{\text{eff}}-1)O^{2}+3s^{2}-2s\right].$ (11)
We now analyze the steady state of the mean-field dynamics. In the stationary regime, the time derivatives vanish, i.e., $\dot{O}=0$, $\dot{s}=0$, and the system settles into fixed points $(O^{*},s^{*})$ of the evolution equations. In the following, we identify these fixed points and discuss separately the disordered solutions with $O^{*}=0$ and the ordered solutions with $|O^{*}|>0$, together with their respective stability properties.
Disordered fixed point
Setting $O^{*}=0$ in Equation˜11 yields
$3(s^{*})^{2}-2s^{*}=0\text{ with solutions }s^{*}=0\text{ or }s^{*}=2/3.$ (12)
The physically relevant disordered state is therefore
$(O^{*},s^{*})=\left(0,\frac{2}{3}\right).$ (13)
Ordered fixed points
For $O^{*}\neq 0$, Equation˜10 requires
$(1-2p_{\text{eff}})-(1-p_{\text{eff}})s^{*}=0\text{ and leading to }s^{*}=\frac{1-2p_{\text{eff}}}{1-p_{\text{eff}}}.$ (14)
Substituting this into Equation˜11 gives the non-zero solutions
$(O^{*})^{2}=\frac{1-4p_{\text{eff}}}{(1-p_{\text{eff}})^{2}},$ (15)
where $p_{\text{eff}}$ can be found from Equation˜9. Hence ordered state exists if $p_{\text{eff}}<1/4$.
III.2.1 Stability of the disordered fixed point
Linearizing Equations˜10 and 11 around the point $(O^{*},s^{*})=\left(0,2/3\right)$, one finds a diagonal Jacobian with eigenvalues
$\lambda_{s}=-1\quad\text{and}\quad\lambda_{O}=\frac{1-4p_{\text{eff}}}{3}.$ (16)
The first eigenvalue is always negative, ensuring stability along the $s$ direction. The second eigenvalue changes sign at $p_{\text{eff}}=1/4$. For $p_{\text{eff}}>1/4$, both eigenvalues are negative and the disordered state is linearly stable, whereas for $p_{\text{eff}}<1/4$, the eigenvalue associated with the order parameter direction becomes positive, making the point a saddle. Thus the disordered fixed point loses stability precisely at $p_{\text{eff}}=1/4$, which identifying the critical line separating the ordered and disordered phases.
Therefore the critical point satisfies
$(1-2c)p_{c}(c)+c=\frac{1}{4},$ (17)
which yields
$p_{c}(c)=\frac{\frac{1}{4}-c}{1-2c},$ (18)
whether this expression falls inside the physical range $p_{c}\in[0;1]$ depends entirely on the value of $c$, and this leads to three distinct regimes:
Regime I ($0<c<1/4$)
In this interval both the numerator and denominator in Equation˜18 are positive, so $p_{c}(c)$ lies strictly between 0 and $1$.
For $p>p_{c}(c)$, one has $p_{\text{eff}}>1/4$, so the eigenvalue associated with the order parameter direction $(\lambda_{O})$ is negative and the disordered fixed point remains stable.
For $p<p_{c}(c)$, the condition $p_{\text{eff}}<1/4$ is reached and the disordered fixed point becomes a saddle. At this point a pair of ordered fixed points with nonzero $\pm O^{*}$ [Equation˜15] emerges and becomes stable. Thus, in the interval $0<c<1/4$, the system displays a genuine order-disorder transition at $p=p_{c}(c)$, with ordered state present whenever $p$ is sufficiently small.
Regime II ($1/4<c<3/4$)
For $c$ in this range, $p_{\text{eff}}$ remains strictly above the threshold $1/4$ for all $p\in[0;1]$. So the disordered fixed point is always stable and no ordered state can be generated. Therefore the whole region $1/4<c<3/4$ corresponds to complete disorder, with no phase transition for any value of $p$.
Regime III ($3/4<c<1$)
When $c$ exceeds $3/4$, in this case Equation˜18 yields a critical point $p_{c}(c)>3/4$.
For $p<p_{c}(c)$, one always has $p_{\text{eff}}>1/4$, ensuring stability of the disordered fixed point.
However, when $p>p_{c}(c)$, the condition $p_{\text{eff}}<1/4$ is reached and the disordered state loses stability, giving rise to two stable ordered solutions $\pm O^{*}$ [Equation˜15]. Thus ordering reappears for $c>3/4$, but now only at sufficiently higher values of $p$, in contrast to Regime I where order appears at smaller values of $p$.
III.2.2 Order parameter scaling near $p_{c}(c)$
In the ordered phase, the non zero stationary order parameter satisfies Equation˜15. To extract the critical behavior of the order parameter, set $\delta=p_{c}(c)-p$. By Taylor expanding $p_{\text{eff}}(p)$ about the critical point $p=p_{c}(c)$, we obtain up-to order $\delta$,
$p_{\text{eff}}(p)\approx p_{\text{eff}}(p_{c})-(1-2c)\delta=\frac{1}{4}-(1-2c)\delta.$ (19)
Substituting Equation˜19 into the numerator of Equation˜15 gives
$1-4p_{\text{eff}}\approx 4(1-2c)\delta$ (20)
and also for the denominator of Equation˜15 we have
$(1-p_{\text{eff}})^{2}\approx\frac{9}{16}+\mathcal{O}(\delta).$ (21)
Using Equations˜20 and 21 in Equation˜15 yields, to leading order in $\delta$
$(O^{*})^{2}\approx\frac{64}{9}(1-2c)\delta+\mathcal{O}(\delta^{2})$ (22)
and therefore
$O^{*}\approx\pm\frac{8}{3}\sqrt{(1-2c)\delta}+\mathcal{O}(\delta^{3/2}).$ (23)
This shows explicitly that for $p<p_{c}(c)$ and near the critical point
$O^{*}\propto(p_{c}(c)-p)^{\frac{1}{2}}.$ (24)
Hence, the order parameter vanishes with exponent $\beta=1/2$, exactly the same as in the original model.
III.2.3 Divergence of the relaxation timescale
To quantify how fast the system returns to the disordered fixed point, we define a relaxation timescale as follows: Consider a small deviation of the order parameter from the disordered fixed point $O^{*}=0$
$O(t)=\epsilon(t)\quad\text{with}\quad\epsilon\ll 1.$ (25)
Linearizing the mean-field equation for $O(t)$ [Equation˜10] around $O^{*}=0$ gives
$\frac{d\epsilon}{dt}=\lambda_{0}\epsilon,\quad\lambda_{0}=\frac{1-4p_{\text{eff}}}{3}.$ (26)
The solution is $\epsilon(t)=\epsilon(0)\exp(\lambda_{0}t)$. Hence, the relaxation timescale is defined as
$\tau=\frac{1}{|\lambda_{0}|}.$ (27)
Near the critical point we insert Equation˜20 into Equation˜26
$\lambda_{0}=\frac{4(2c-1)\delta}{3},\quad\delta=p-p_{c}(c).$ (28)
Hence, up to a finite prefactor
$\lambda_{0}\propto\delta,\text{ yielding }\tau\propto\delta^{-1}\propto(p-p_{c}(c))^{-1}.$ (29)
Therefore, the relaxation timescale diverges at the critical point with an effective exponent $\nu z=1$, exactly in the same way as in the original model.
III.3 Mean-field calculations for nonconformists (5)
We now introduce the nonconformists in the second version according to Equation˜5. Following we define three basic probabilities
$P_{+}=\operatorname{\mathrm{Prob}}(\mu_{ij}o_{j}=+1)=(1-p)f_{+}+pf_{-},$ (30)
$P_{-}=\operatorname{\mathrm{Prob}}(\mu_{ij}o_{j}=-1)=(1-p)f_{-}+pf_{+}.$ (31)
$P_{0}=\operatorname{\mathrm{Prob}}(\mu_{ij}o_{j}=0)=f_{0}.$ (32)
From Equation˜5, the mean-field transition rates between the three opinion states can be written in terms of $P_{+}$ and $P_{-}$ as
$\omega_{+1\to-1}=c(P_{+}+f_{0}),\quad\omega_{-1\to+1}=c(P_{-}+f_{0}),$ (33)
$\omega_{-1\to 0}=P_{+},\quad\omega_{0\to+1}=(1-c)P_{+}+cP_{-},$ (34)
$\omega_{0\to-1}=(1-c)P_{-}+cP_{+},\quad\omega_{+1\to 0}=P_{-}.$ (35)
One interesting aspect of this model is that it allows extreme switches, i.e., transitions from opinion $1\to-1$ or vice versa, which the original BChS dynamics does not permit but can be incorporated in a different way [61].
Using the transition rates listed above, the mean-field time evolution of the fractions of positive and negative agents follows directly from the corresponding gain-loss balances. In particular the rate equation for $f_{+}$ and $f_{-}$ take the form
$\begin{split}\frac{df_{+}(t)}{dt}=f_{0}[(1-c)P_{+}+cP_{-}]\\ +f_{-}[c(P_{-}+f_{0})]-f_{+}[c(P_{+}+f_{0})+P_{-}],\end{split}$ (36)
$\begin{split}\frac{df_{-}(t)}{dt}=f_{0}[(1-c)P_{-}+cP_{+}]\\ +f_{+}[c(P_{+}+f_{0})]-f_{-}[c(P_{-}+f_{0})+P_{+}].\end{split}$ (37)
Adding and subtracting Equations˜36 and 37 gives the closed evolution equations for the order parameter $O$ and activity $s$. However, the activity equation is the same as in the original mean-field model:
$\frac{ds(t)}{dt}=-\frac{1}{2}\left[(2p-1)O^{2}+3s^{2}-2s\right];$ (38)
$\frac{dO(t)}{dt}=[((1-2c)(1-2p)-2c)(1-s)-(p+2c(1-p))s]O(t).$ (39)
A steady state is obtained from $\dot{O}(t)=0$ and $\dot{s}(t)=0$.
Disordered fixed point
Setting $O^{*}=0$ in Equation˜38 yields the solution $s^{*}=0$ or $s^{*}=2/3$. The physically relevant disordered fixed point is therefore $(O^{*},s^{*})=(0,2/3)$, exactly same as in the original model as well as in the model of nonconformist of first version [Equation˜4].
Ordered fixed points
For $O^{*}\neq 0$, the factor multiplying $O$ in $\dot{O}$ in Equation˜39 must vanish, giving an equation for
$s^{*}=\frac{(1-2c)(1-2p)-2c}{(1-2c)(1-p)}.$ (40)
Substituting this $s^{*}$ into the steady state condition $\dot{s}=0$ in Equation˜38 with $O^{*}\neq 0$ yields the ordered branch in the standard form
$(O^{*})^{2}=\frac{2s^{*}-3(s^{*})^{2}}{2p-1}.$ (41)
III.3.1 Stability of the disordered fixed point
Linearizing Equations˜38 and 39 around $(O^{*},s^{*})=(0,2/3)$, the Jacobian is diagonal and the eigenvalue controlling the order parameter direction is
$\lambda_{O}=\frac{1-4p-8c+8cp}{3},$ (42)
while the $s$ direction eigenvalue is $\lambda_{s}=-1<0$. Therefore, the disordered fixed point is stable when $\lambda_{O}<0$ (with both eigenvalues negative) and becomes a saddle when $\lambda_{O}>0$.
The requirement that an ordered branch exists imposes the condition that the steady state activity satisfies $s^{*}>2/3$, for which the rhs in Equation˜41 remains positive. Using the explicit form of $s^{*}$ in Equation˜40, one verifies that $s^{*}>2/3$ holds exactly when $\lambda_{O}>0$. So, this is precisely the condition for the emergence of order.
The transition line is obtained from $\lambda_{O}=0$, i.e.,
$p_{c}(c)=\frac{8c-1}{4(2c-1)},\quad c\neq\frac{1}{2}.$ (43)
Following we discuss two distinct regimes:
Regime I ($0<c<1/8$)
For $c$ in this particular regime, the expression of critical point in Equation˜43 yields a physical value i.e. $p_{c}(c)\in[0,1]$ and there exists an interval $0\leq p\leq p_{c}(c)$ where $\lambda_{O}>0$. Therefore in this region, the disordered fixed point is a saddle and two stable ordered fixed points appear. Thus in the interval $0<c<1/8$, the system displays a genuine order-disorder transition at $p=p_{c}(c)$.
Regime II ($c>1/8$)
For $c>1/8$, for all value of $p\in[0,1]$, $\lambda_{O}<0$ and the system remains disordered for all $p$.
III.3.2 Order parameter scaling near $p_{c}(c)$
In the ordered phase, the non-zero stationary order parameter satisfies Equation˜41. To extract the critical behavior of the order parameter we introduce $\delta=p_{c}(c)-p$—where $p_{c}(c)$ can be found from Equation˜43— and expand all quantities for small $\delta$.
Near the transition, the activity $s^{*}$ in Equation˜40 satisfies
$s^{*}\approx\frac{2}{3}+\frac{16}{9}(1-2c)\delta+\mathcal{O}(\delta^{2}).$ (44)
This is obtained by Taylor expanding the exact expression for $s^{*}$ around $p=p_{c}(c)$.
Substituting Equation˜44 into the numerator of Equation˜41 gives
$2s^{*}-3(s^{*})^{2}\approx\frac{32}{9}(2c-1)\delta+\mathcal{O}(\delta^{2}).$ (45)
For the denominator of Equation˜41 we expand
$2p-1\approx(2p_{c}(c)-1)-2\delta.$ (46)
Using Equations˜45 and 46 in Equation˜41 yields, to leading order in $\delta$
$(O^{*})^{2}\approx\frac{32}{9}\frac{(2c-1)}{(2p_{c}(c)-1)}\delta+\mathcal{O}(\delta^{2}).$ (47)
Taking the square root gives the scaling form
$O^{*}\approx\sqrt{\frac{32}{9}\frac{(2c-1)}{(2p_{c}(c)-1)}}\delta^{\frac{1}{2}}.$ (48)
Since, the ordered solution $O^{*}$ exists only for $c<1/8$ and $p<p_{c}(c)<1/4$, the quantity appearing inside the square root in Equation˜48 is always positive. Thus, for $p<p_{c}(c)$ and sufficiently close to the critical point,
$O^{*}\propto(p_{c}(c)-p)^{\frac{1}{2}}$ (49)
showing explicitly that the order parameter vanishes with exponent $\beta=1/2$, exactly the same mean-field value as in the original model.
III.3.3 Divergence of the relaxation timescale
The divergence of the relaxation timescale can be obtained by following exactly the same steps as in the earlier analysis for nonconformists [Equation˜4] which is given in Section˜III.2.3. Linearizing, the order parameter dynamics [Equation˜39] around the disordered fixed point $(O^{*},s^{*})=(0,2/3)$ leads to an exponential relaxation controlled by the eigenvalue $\lambda_{O}$ given in Equation˜42.
Now, we define $\delta=p-p_{c}(c)$ and expand $\lambda_{O}(p,c)$ about $p=p_{c}(c)$. Since, $\lambda(p_{c}(c),c)=0$ by definition of the critical line, the leading term is linear in $\delta$. Therefore, we obtain, to leadimg order
$\lambda_{O}(p,c)\approx\frac{4(2c-1)}{3}\delta+\mathcal{O}(\delta^{2}).$ (50)
Thus up to a finite prefactor
$\lambda_{O}\propto\delta\text{ leading to }\tau\propto\delta^{-1}\propto(p-p_{c}(c))^{-1}.$ (51)
Thus, the relaxation time diverges linearly at the critical point with the same effective exponent $\nu z=1$ as in the original mean-field model.
III.4 Mean-field calculations for inflexibles
III.4.1 Inflexibles adhered to extreme opinions
We now consider a variant of the original kinetic exchange model in which inflexibles are restricted to agents holding extreme opinions as discussed for intransigents and the quenched case in Section III B in Reference 52. A fraction $z_{+}$ of agents with opinion +1+1 and a fraction $z_{-}$ of agents with opinion 1-1 are inflexibles. Neutral agents with zero opinion are always flexible.
Since, only extreme agents can be inflexibles, the transition rates out of $\pm 1$ states are reduced. In particular,
$\omega_{+1\to 0}=(1-z_{+})P_{-},\qquad\omega_{-1\to 0}=(1-z_{-})P_{+},$ (52)
$\omega_{+1\to-1}=0,\,\omega_{-1\to+1}=0,\,\omega_{0\to+1}=P_{+},\,\omega_{0\to-1}=P_{-},$ (53)
where probabilites $P_{+}$ and $P_{-}$ are again given by Equations˜30 and 31. Using the transition rates listed above, the rate equation for $f_{+}$ and $f_{-}$ takes the form
$\frac{df_{+}(t)}{dt}=P_{+}f_{0}-(1-z_{+})P_{-}f_{+},$ (54)
$\frac{df_{-}(t)}{dt}=P_{-}f_{0}-(1-z_{-})P_{+}f_{-}.$ (55)
Now, using
$\begin{split}P_{-}f_{+}-P_{+}f_{-}=psO,\\ P_{-}f_{+}+P_{+}f_{-}=\left[\frac{s^{2}}{2}-\frac{(1-2p)}{2}O^{2}\right],\end{split}$ (56)
$P_{+}-P_{-}=(1-2p)O,\quad P_{+}+P_{-}=s,$ (57)
$z_{+}P_{-}f_{+}=\frac{z_{+}s^{2}}{4}+\frac{z_{+}psO}{2}+\frac{z_{+}O^{2}}{2}\left(p-\frac{1}{2}\right),$ (58)
$z_{-}P_{+}f_{-}=\frac{z_{-}s^{2}}{4}-\frac{z_{-}psO}{2}+\frac{z_{-}O^{2}}{2}\left(p-\frac{1}{2}\right)$ (59)
and introducing the ratio $\rho=z_{+}/z_{-}$ so that
$z_{+}+z_{-}=z,\quad z_{+}-z_{-}=\left(\frac{\rho-1}{\rho+1}\right)z.$ (60)
By subtracting and adding Equations˜54 and 55, one obtains the closed evolution equation for the order parameter $O(t)$ and activity $s(t)$ as
$\begin{split}\frac{dO(t)}{dt}=(1-s)(1-2p)O-psO+\left(\frac{\rho-1}{\rho+1}\right)\frac{zs^{2}}{4}\\ +\frac{zpsO}{2}+\left(\frac{2p-1}{4}\right)\left(\frac{\rho-1}{\rho+1}\right)zO^{2},\end{split}$ (61)
$\begin{split}\frac{ds(t)}{dt}=s\left[1+\left(\frac{z}{4}-\frac{3}{2}\right)s\right]+\frac{pz}{2}\left(\frac{\rho-1}{\rho+1}\right)sO\\ +\frac{(1-2p)}{2}\left(1-\frac{z}{2}\right)O^{2}.\end{split}$ (62)
Following we discuss three distinct cases:
Case I: Symmetric inflexibles $(\rho=1)$
When $\rho=1$, we have $z_{+}=z_{-}$ and therefore $z_{+}-z_{-}=0$. As a result, Equations˜61 and 62 reduce to
$\frac{dO(t)}{dt}=\left[(1-s)(1-2p)-\left(1-\frac{z}{2}\right)ps\right]O,$ (63)
$\frac{ds(t)}{dt}=s\left[1+\left(\frac{z}{4}-\frac{3}{2}\right)s\right]+\frac{(1-2p)}{2}\left(1-\frac{z}{2}\right)O^{2}.$ (64)
One can find the disordered fixed point by setting $O^{*}=0$ in Equation˜64, which yields the solution $s^{*}=4/(6-z)$. The physically relevant disordered fixed point is therefore $O^{*},s^{*}]=[0,4/(6-z)]$. Linearizing Equation˜63 around $O^{*},s^{*}]=[0,4/(6-z)]$, the eigenvalue controlling the order parameter direction is
$\lambda_{O}=\left(\frac{2-z}{6-z}\right)(1-4p).$ (65)
The critical point is obtained from $\lambda_{O}=0$, i.e.,
$p_{c}=\frac{1}{4},$ (66)
which is exactly same as in the original model.
Case II: Positive imbalance ($\rho>1$)
Now we consider a small deviation from the symmetric case by introducing a positive imbalance between the two types of inflexibles. We set
$\rho=1+\epsilon,\qquad 0<\epsilon\ll 1,$ (67)
so that $z_{+}-z_{-}>0$.
Substituting this into Equation˜61 and evaluating it at $O^{*}=0$, we obtain, to linear order in $\epsilon$,
$\frac{dO(t)}{dt}\Big|_{O^{*}=0}=\frac{z(s^{*})^{2}}{8}\epsilon.$ (68)
This contribution is strictly positive for any finite inflexible fraction $z>0$. Hence, irrespective of how small the imbalance $\epsilon$ is, the disordered state $O^{*}=0$ is no longer a fixed point of the dynamics. Therefore, the notion of the disordered phase ceases to exist once $\rho$ deviates slightly from unity along the positive side. The order parameter remains finite for all values of $p$, and its magnitude increases with increasing inflexible fraction. So, a positive inflexible imbalance acts as a positive bias leading to the suppression of order disorder phase transition.
Case III: Negative imbalance ($\rho<1$)
An analogous conclusion follows for a negative imbalance. Setting
$\rho=1-\epsilon,\qquad 0<\epsilon\ll 1$ (69)
we have $z_{+}-z_{-}<0$.
Substituting this into Equation˜61 and evaluating it at $O^{*}=0$, we obtain, to linear order in $\epsilon$,
$\frac{dO(t)}{dt}\Big|_{O^{*}=0}=-\frac{z(s^{*})^{2}}{8}\epsilon.$ (70)
This contribution is strictly negative for any finite inflexible fraction $z>0$. Hence, irrespective of how small the imbalance $\epsilon$ is, the disordered state $O^{*}=0$ is again no longer a fixed point of the dynamics. Consequently, the notion of the disordered phase ceases to exist once $\rho$ deviates slightly from unity along the negative side. The order parameter remains finite for all values of $p$ with a negative sign, and its magnitude increases with increasing inflexible fraction. Thus, a negative inflexible imbalance acts as a negative bias, leading again to the suppression of the order disorder transition.
III.4.2 Inflexibles adhered to specific opinion
We now consider another version of inflexible, where intransigence is restricted to actors holding a specific opinion state, as discussed the intransigents and the quenched case in Section III C in Reference 52. Here, however, we analyze the annealed mean-field version of this model.
Inflexibles restricted to opinion +1+1
We first take a fraction $z_{+}=z$ of agents with opinion +1+1 to be inflexibles, while $z_{-}=z_{0}=0$. In terms of the imbalance parameter, this corresponds to the extreme limit $\rho=z_{+}/z_{-}\to\infty$, which is the boundary case of a positive imbalance. Based on the previous analysis, we therefore expect behavior analogous to that induced by a positive bias.
In this case, only the transition $\omega_{+1\to 0}$ is affected, being suppressed by a factor of (1z)(1-z). The resulting evolution equation for the order parameter becomes,
$\begin{split}\frac{dO(t)}{dt}=\left[(1-s)(1-2p)+\frac{ps}{2}(z-2)\right]O+\frac{zs^{2}}{4}\\ +\frac{z(2p-1)}{4}O^{2}.\end{split}$ (71)
Evaluating this equation at $O^{*}=0$ yields
$\frac{dO(t)}{dt}\Big|_{O^{*}=0}=\frac{z(s^{*})^{2}}{4}.$ (72)
Since, the contribution is strictly positive for any finite inflexible fraction $z>0$, the disordered fixed point $O^{*}=0$ ceases to exist. Consequently, the order disorder phase transition is suppressed, as inflexibles restricted to the +1+1 state act as an effective positive bias.
Inflexibles restricted to opinion 1-1
Next, we consider the opposite extreme case with $z_{-}=z$ and $z_{+}=z_{0}=0$, corresponding to $\rho=0$. This represents the boundary case of a negative imbalance, and we therefore expect behavior analogous to that induced by a negative bias.
Here, only the transition $\omega_{-1\to 0}$ is modified. The resulting order parameter evolution equation reads
$\begin{split}\frac{dO(t)}{dt}=\left[(1-s)(1-2p)+\frac{ps}{2}(z-2)\right]O\\ -\frac{zs^{2}}{4}-\frac{z(2p-1)}{4}O^{2}.\end{split}$ (73)
Evaluating this equation at $O^{*}=0$ yields
$\frac{dO(t)}{dt}\Big|_{O^{*}=0}=-\frac{z(s^{*})^{2}}{4}.$ (74)
Since, the contribution is strictly negative for any finite inflexible fraction $z>0$, the disordered fixed point $O^{*}=0$ again ceases to exist in this case. Consequently, the order-disorder phase transition is suppressed, as inflexibles restricted to the 1-1 state act as an effective negative bias.
Inflexibles restricted to opinion 0
We now consider the case where inflexibles are restricted exclusively to the neutral state i.e. $z_{+}=z_{-}=0$ and $z_{0}=z$. Since, there is no imbalance between the extreme opinions, this corresponds to a zero bias situation, and therefore one expects the system to behave similarly to the unbiased case, with an order disorder phase transition at a finite critical point.
In this case, only the transitions $0\to+1$ and $0\to-1$ are affected by the presence of inflexibles. The resulting mean-field evolution equations for the order parameter $O$ and the activity $s$ reads,
$\frac{dO(t)}{dt}=\left[(1-s)(1-2p)(1-z)-ps\right]O,$ (75)
$\frac{ds(t)}{dt}=s(1-s)(1-z)-\frac{s^{2}}{2}+\frac{(1-2p)}{2}O^{2}.$ (76)
Setting $O^{*}=0$ in Equation˜76, one finds the physically relevant disordered fixed point
$(O^{*},s^{*})=\left(0,\frac{2(1-z)}{3-2z}\right).$ (77)
Linearizing Equation˜75 around this disordered fixed point yields the eigenvalue associated with the order-parameter direction,
$\lambda_{O}=\left(\frac{1-z}{3-2z}\right)(1-4p).$ (78)
The critical point is therefore given by
$p_{c}=\frac{1}{4},$ (79)
which is exactly the same as in the original model and coincides with the result reported in Section C of Reference 52 for the quenched version of neutral intransigents.
III.4.3 Inflexibles chosen independently of their initial opinion
We now turn to another version of inflexibles, where a fraction $z$ of agents is chosen uniformly at random, independently of their initial opinion, as discussed in Section III A of Reference 52 for intransigent and the quenched case. Here, we analyze the corresponding annealed mean-field version.
Since inflexibles never update their opinions, all transition rates involving opinion updates are suppressed by a factor (1z)(1-z). With this modification, the mean-field evolution equation for the order parameter becomes
$\frac{dO(t)}{dt}=(1-z)\left[(1-s)(1-2p)-ps\right]O,$ (80)
while the activity evolves according to
$\frac{ds(t)}{dt}=(1-z)\left[s(1-s)-\frac{s^{2}}{2}+\frac{(1-2p)}{2}O^{2}\right].$ (81)
Setting $O^{*}=0$, Equation˜81 yields the same disordered fixed point as in the original model, $(O^{*},s^{*})=(0,2/3)$, independent of the inflexible fraction $z$.
Linearizing Equation˜80 around this disordered fixed point gives the eigenvalue
$\lambda_{O}=\frac{1-z}{3}(1-4p).$ (82)
The critical point is therefore
$p_{c}=\frac{1}{4},$ (83)
which is unchanged by the presence of annealed inflexibles. This is in clear contrast to the quenched case reported in Equation (11) of Section III A of Reference 52, where the critical point shifts with the inflexible fraction.
Substituting $O^{*}\neq 0$ and the stationary activity $s^{*}=2/3$ into Equation˜81 yields the ordered fixed points for $p<1/4$
$(O^{*})^{2}=\frac{1-4p}{(1-p)^{2}},$ (84)
which is independent of the inflexible fraction $z$. Hence in the annealed case, inflexibles merely rescale the timescale of the dynamics but do not affect the location of the critical point or the stationary order parameter.
III.5 Monte Carlo simulations
The simulation is performed on the fully connected graph with $N$ nodes. Symbols $\{\cdots\}$ and $\langle\cdots\rangle$ stand for a time average over the last $\tau$ time steps and average over $R$ independent simulations, respectively. All simulations are performed for a system size $N=1024$ and averaged over $R=10^{3}$ independent simulations. For a single simulation, the results are gathered from the last $\tau=0.2T$ time steps and $T=10^{3}$. Both, quenched and annealed disorder are considered. For the annealed disorder, every selected actor decides to behave as nonconformist (inflexible) with the probability $c$ ($z$). For the quenched disorder, the positions of nonconformists (inflexibles) in a single simulation are fixed on the graph and their fraction is $c$ ($z$).
Figure 1: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜4 (upper panel) and Equation˜5 (lower panel). The annealed disorder is assumed. $N=1024$, $T=10^{3}$, $R=10^{3}$, $\tau=0.2T$
III.5.1 Discrete opinions
For discrete opinions, we have $o_{i}\in\{-1,0,+1\}$ and $\mu_{ij}\pm 1$. The fraction of negative values of $\mu_{ij}$ is $p$.
In Figure˜1 the mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜4 (upper panel) and Equation˜5 (lower panel) are presented. For both variants [Equations˜4 and 5] the annealed disorder is assumed, that is, the selected actor applies the rule (4) or (5) instead of the rule (1) with probability $c$.
Figure˜2 shows the mean order parameter $\langle\{O\}\rangle$ for various fractions $z$ of inflexibles with discrete opinions. The quenched (upper panel) or annealed (lower panel) disorder is assumed. For the quenched disorder fraction of $z$ actors, marked as inflexibles, has the fixed opinion +1+1. For the annealed disorder, the selected actor, with probability $z$ takes opinion +1+1 instead of following the rule (1).
Figure 2: Mean order parameter $\langle\{O\}\rangle$ for various fractions $z$ of inflexibles with discrete opinions. The inflexibles are adhered to opinion +1+1. The quenched (upper panel) or annealed (lower panel) disorder is assumed. $N=1024$, $T=10^{3}$, $R=10^{3}$, $\tau=0.2T$
III.5.2 Continuous opinions
For continuous opinions, we have $o_{i}\in[-1;+1]$ and also $\mu_{ij}\in[-1;+1]$. The fraction $c$ of agents is considered nonconformist and $p$ is a fraction of negative values of $\mu_{ij}$.
Figure˜3 shows the order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with continuous opinions. In the upper (bottom) panel the results for nonconformists introduced according to Equation˜4 (Equation˜5) are presented.
Figure 3: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists for continuous opinions. The upper panel corresponds to nonconformists introduced according to Equation˜4, while the lower shows results for implementing rule given by Equation˜5. The quenched disorder is assumed. $N=1024$, $T=10^{3}$, $R=10^{3}$, $\tau=0.2T$
In Figure˜4 the order parameter $\langle\{O\}\rangle$ is presented for various fractions $z$ of inflexibles with continuous opinions. Inflexibles are adhered to opinion +1+1. For continuous opinions only quenched disorder is available.
Figure 4: Mean order parameter $\langle\{O\}\rangle$ for various fractions $z$ of inflexibles for continuous opinions. The inflexibles are adhered to opinion +1+1. Only the quenched disorder may be considered. $N=1024$, $T=10^{3}$, $R=10^{3}$, $\tau=0.2T$
Figure 5: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜4. The solid line represent the closed form expression of the order parameter given in Equation˜15 while the points denote the simulation results. The analytical prediction shows excellent agreement with the simulation data. All simulations were done for the system size $N=1024$
Figure 6: Mean order parameter $\langle\{O\}\rangle$ for various fractions $c$ of nonconformists with discrete opinions modeled according to Equation˜5. The solid line represent the closed form expression of the order parameter given in Equation˜41 while the points denote the simulation results. The analytical prediction shows excellent agreement with the simulation data. All simulations were done for the system size $N=1024$
IV Discussion
Figure 7: Mean order parameter $\langle\{O\}\rangle$ for various fractions $z$ of inflexibles chosen independently of their initial opinion (Section˜III.4.3). Discrete opinions and annealed disorder are assumed. The solid line represents the closed form expression of the order parameter given in Equation˜84 while the points denote the simulation results. All simulations were done for the system size $N=1024$
For nonconformists introduced according to Equation˜4 the mean-field approach leads to conclusions that an ordered phase is available when the fraction of nonconformists is $c<1/4$ or $c>3/4$. The mean-field calculations applied to the nonconformists introduced according to Equation˜5 show that the ordered state may be reached only when $c<1/8$.
Figure 8: Mean order parameter $\langle\{O\}\rangle$ for fractions $z=0.2$ (left panel), $z=0.5$ (middle panel) and $z=0.8$ (right panel) of inflexibles and discrete/continuous opinions and the quenched/annealed disorder. The inflexibles keep opinion (+1+1) with probability $z$ (for annealed disorder) or have opinion (+1+1) and fixed position on network (for quenched disorder). In the latter the fraction $z$ of nodes is marked as occupied by inflexibles. $N=1024$, $T=10^{3}$, $R=10^{3}$, $\tau=0.2T$
The simulation results for discrete opinions and annealed disorder agree perfectly with the mean-filed theory for nonconformists introduced accordingly to Equation˜4 and Equation˜5 as presented in Figures˜5 and 6, respectively.
For inflexibles adhered to extreme opinions (Section˜III.4.1), the symmetric point $\rho=1$ (with the fractions of inflexibles adhered to opinion +1+1 equal to the fractions of inflexibles adhered to opinion 1-1) is special. It is the only value of $\rho$ for which the order parameter equation is invariant under $O\to-O$ and admits a disordered fixed point. Any infinitesimal deviation from $\rho=1$ destroys this symmetry and suppresses the phase transition. In this sense, $\rho=1$ is a structurally unstable point protected by symmetry that separates unbiased dynamics from biased ones.
The introduction of inflexibles (Section˜III.4.2) adhered to a specific opinion (for example, +1+1) leads to a subsequent increase in the order parameter with the fraction $z$ of them (up-to $O\equiv 1$, independently on $p$, for $z=1$). In such a case, the order disorder phase transition is suppressed (see Figure˜4).
For inflexibles chosen independently of their initial opinion (Section˜III.4.3) the mean order parameter $\langle\{O\}\rangle$ is independent of their fraction $z$ as predicted by mean-field calculations (84) as presented in Figure˜7.
The results of simulations are qualitatively the same independently on assumed discrete (Figures˜1 and 2) or continuous (Figures˜3 and 4) opinions. Rather surprisingly, for nonconformists introduced via Equation˜4, Monte Carlo simulations for quenched disorder (when the positions of nonconformists are fixed during a given simulation) produce a little better match with the mean-filed calculations. On the other hand, simulations with quenched disorder are better motivated from a sociological perspective—then actors are nonconformists (or inflexibles) and not only behave as nonconformists (or inflexibles). The results for continuous opinions (Figures˜3 and 4) are provided for the assumed quenched disorder.
Figure˜8 shows the mean order parameter $\langle\{O\}\rangle$ for fractions $z=0.2$ (Figure˜8, left panel), $z=0.5$ (Figure˜8, middle panel) and $z=0.8$ (Figure˜8, right panel) of inflexibles and discrete/continuous opinions and the quenched/annealed disorder. For continuous opinions, only quenched disorder is possible to simulate. Independently of the fraction $z$ of inflexibles, the variant with quenched disorder outperforms the variant with annealed disorder to reach unanimity of opinions (i.e., increasing $\langle\{O\}\rangle$). In version with annealed disorder, every actor my suddenly (with probability 1z1-z) leave his/her opinion (for example +1+1), while for quenched disorder (with fixed positions of inflexibles) inflexibles extreme opinions never change, and thus reaching higher values of $\langle\{O\}\rangle$ is easier.
In the model of Crokidakis et al. [52], contrarian like behavior originates from negative interactions: during such interactions, an agent is forced to move away from the neighbor’s opinion, although the same agent may behave normally in other interactions. Thus, opposition there is interaction-driven and temporary. In contrast, here, for both versions, nonconformist behavior is an intrinsic agent property, a nonconformist systematically reverses the response of a normal agent and opposes the neighbor regardless of whether the interaction is positive or negative, which is more realistic. So, the bottom line is that the contrarian behavior in Reference 52 is completely different from the behavior of nonconformists presented in current research. Moreover, while Crokidakis et al. study intransigents in several possible ways, their analysis is primarily numerical, and here the results of Monte Carlo simulation for inflexibles are supported with mean-field calculations.
In conclusion, in this paper, the BChS model is extended by introducing nonconformists and inflexibles. With direct Monte Carlo simulations and mean-field calculations we show the influence of fraction $c$ of nonconformists and fraction $z$ of inflexibles on the order parameter $O$, which captures net polarization. With the mean-field calculations, we identified fractions $c$ of nonconformists that lead to ordered phases. For Monte Carlo simulations, discrete/continuous opinions were introduced, and annealed/quenched disorder was considered. The qualitative results of introducing nonconformists (inflexibles) in various ways (discrete/continuous opinions and annealed/quenched disorder) are roughly the same. However, for the BChS model extended by inflexibles adhered to opinion +1+1, we can see systematic shift of the mean order parameter to its higher values for quenched disorder than for the variant with annealed disorder.
Acknowledgements.
KM thanks PS for her hospitality in the University of Calcutta, Kolkata. KM visit to Kolkata was supported by Excellence Initiative—Research University (IDUB) in AGH University of Kraków. This research was supported by a subsidy from the Polish Ministry of Science and Higher Education.